# Adaptive Smooth Non-Stationary Bandits

### v1

@@ Appendix B Proof of Dynamic Regret Upper Bound (Theorem 5)
The main idea is that on each significant phase $[\tau_{i},\tau_{i+1},$), any arm $a$ which is at one point optimal, i.e. $\delta_{t}(a)=0$ for some $t\in[\tau_{i},\tau_{i+1})$, must have a gap function $t\mapsto\delta_{t}(a)$ with large variation within the phase since the gap must also at some point be large because of our notion of significant regret (see Fact 16). More generally, if the gap function is $\left\lfloor\beta\right\rfloor$ times differentiable, then we can find an order $\left\lfloor\beta\right\rfloor$ critical point of the gap function across $\left\lfloor\beta\right\rfloor$ different phases using Rolle’s Theorem. Using the definition of Hölder function (Definition 1), we can then bound the derivatives of the gap functions using this critical point. Such bounds can in turn be plugged into an order-$\left\lfloor\beta\right\rfloor$ Taylor approximation of the gap function. Ultimately, these calculations allow us to relate the phase length $\tau_{i+1}-\tau_{i}$ to the smoothness parameters $\beta,\lambda$. The key claim is that each phase must be roughly at least length $T^{\frac{2\beta}{2\beta+1}}\lambda^{-\frac{2}{2\beta+1}}K^{\frac{1}{2\beta+1}}$ which gives the desired regret bound.
@@
$\sum_{i=0}^{{\tilde{L}}}\sqrt{K\cdot(\tau_{i+1}-\tau_{i})}=\sqrt{KT}+\sum_{i:\tau_{i+1}-\tau_{i}>(m+1)\cdot K}\sqrt{K\cdot(\tau_{i+1}-\tau_{i})}+\sum_{i:\tau_{i+1}-\tau_{i}\leq(m+1)\cdot K}\sqrt{K\cdot(\tau_{i+1}-\tau_{i})},$
@@
Note that, going forward in the proof, we’ll constrain our attention to significant phases $[\tau_{i},\tau_{i+1})$ such that $\tau_{i+1}\neq T+1$ as the $\sqrt{KT}$ term on the above RHS accounts for this regret contributed by this final phase.
@@
First, suppose $m=0$. Then, we have by Fact 16 that $F(x)=\delta_{x\cdot T}(a)\geq\sqrt{\frac{K}{\tau_{i+1}-\tau_{i}+1}}$ for some $x\in[\tau_{i}/T,\tau_{i+1}/T]$. Combining this with our above display, there exists $x\in[\tau_{i}/T,\tau_{i+1}/T]$ such that for $m=0$:
@@
Then, taking an order-$(m-1)$ Taylor expansion with Lagrange remainder of $F$ about a root $x_{1}\in[\tau_{i}/T,\tau_{i+1}/T)$ (which we already argued exists since $a$ is optimal at some round in $[\tau_{i},\tau_{i+1})$), we have there exists $\xi\in[\tau_{i}/T,\tau_{i+1}/T)$ such that:
$\forall x\in\bigg{[}\frac{\tau_{i}}{T},\frac{\tau_{i+1}}{T}\bigg{)}:|F(x)|$ $=|F(x)-F(x_{1})|$ ($F(x_{1})=0$)
@@
We next analyze the short significant phases $[\tau_{i},\tau_{i+1})$ where $\tau_{i+1}-\tau_{i}\leq(m+1)\cdot K$. The difficulty here is that we cannot directly apply the same argument as we did for long phases since there may not exist $m+1$ different rounds where an arm is optimal within the phase. To get around this, we’ll concatenate different short phases together and construct pseudo phases where we can apply the argument as we made for long phases.
@@
We’ll now further decompose the dynamic regret over short phases using Jensen’s inequality and the fact that each pseudo phase $[n_{j},n_{j+1})$ can contain at most $(m+1)$ short phases $[\tau_{i},\tau_{i+1})$ since all phases are length at least $K$ (Fact 15):
@@
where $J_{1}$ and $J_{2}$ are respectively the number of filled and unfilled pseudo phases. We’ll first bound $J_{1}$.
@@ C.1   Proof of Theorem 9
$\delta_{\ell}(a)\sum_{t\in P_{\ell}\cap[1,t_{a}-1]}\frac{1}{|\mathcal{S}_{t}|}\leq C_{2}^{2}\frac{\log(T)\sum_{t\in P_{\ell}\cap[1,t_{a}-1]}|\mathcal{S}_{t}|^{-1}}{\delta_{\ell}(a)\sum_{t\in P_{\ell}\cap[1,t_{a}-1]}|\mathcal{S}_{t}|^{-1}}=\frac{C_{2}^{2}\log(T)}{\delta_{\ell}(a)}.$
@@ C.2   Proof of Theorem 11
$\leq C_{2}\sum_{i=1}^{K}\sum_{a\in[K]}\boldsymbol{1}\{a\in\mathcal{S}_{t_{i}-1}\backslash\mathcal{S}_{t_{i}}\}\sqrt{\sum_{t=1}^{t_{i}-1}\frac{\log(T)}{|\mathcal{S}_{t}|}}$ (from Equation 4)
$\leq C_{2}\sqrt{K\sum_{i=1}^{K}\sum_{t=1}^{t_{i}-1}\frac{\log(T)}{|\mathcal{S}_{t}|}}$ (Jensen’s inequality)
$\leq C_{2}\sqrt{K\sum_{i=1}^{K}|\mathcal{S}_{t_{i-1}}|\sum_{t=t_{i-1}}^{t_{i}}\frac{\log(T)}{|\mathcal{S}_{t_{i-1}}|}}$ (from $\mathcal{S}_{t}\supseteq\mathcal{S}_{t+1}$)
$\leq C_{2}\sqrt{KT\log(T)}.$
@@ C.3   Proof of Theorem 12
$\sum_{s=s_{1}}^{s_{2}}\frac{1}{|\mathcal{A}_{s}|}\geq\max_{a^{\prime}\in[K]}\sum_{s=s_{1}}^{s_{2}}\frac{\hat{\delta}_{s}(a^{\prime},{a^{\sharp}})}{|\mathcal{A}_{s}|}\geq C_{5}\sqrt{\sum_{s=s_{1}}^{s_{2}}\frac{\log(T)}{|\mathcal{A}_{s}|}}\implies\sum_{s=s_{1}}^{s_{2}}\frac{1}{|\mathcal{A}_{s}|}\geq C_{5}^{2}\log(T).$
In particular, this means for $C_{5}>1$ in Algorithm 1: we will have
@@
which violates the definition of the safe arm (Definition 10) for $C_{3}=1$ for $C_{5}$ sufficiently large. Thus, we conclude that ${a^{\sharp}}\in\cap_{t=1}^{T}\mathcal{A}_{t}$.
@@
The first bound is order $\sqrt{\sum_{s=s_{1}}^{s_{2}}\frac{\log(T)}{|\mathcal{A}_{s}|}}$ via the definition of ${a^{\sharp}}$ and the second sum is also the same order by our concentration estimate Equation 13 and eviction criterion (Algorithm 1). This establishes Equation 4 for appropriately large $C_{2}$ (in terms of $C_{5}$). Thus, $\{\hat{t}_{i}\}_{i\in[K]}$ are valid eviction times w.r.t. safe armsets $\{\mathcal{A}_{t}\}_{t=1}^{T}$.
@@ C.4   Proof of Theorem 13
Fix a positive integer $T$, number of arms $K\in[2,T]\cap\mathbb{N}$ and a real number $\Delta\in[0,\sqrt{K/T}]$. Let $\mathscr{E}^{\prime}$ be the class of all environments such that $\sum_{t=1}^{T}\mathbb{E}_{a\sim\operatorname{\operatorname{Unif}}\{[K]\}}[\delta_{t}(a)]\leq R$, and such that $T$ is a valid eviction time w.r.t. initial time $1$ and threshold $C_{2}=1$ in Equation 4. Then, for any algorithm $\pi$, we have:
@@
As in the proof of Theorem 3, we’ll follow a Le Cam’s method style of argument for showing minimax lower bounds in stationary bandits (e.g. Lattimore and Szepesvári, 2020, Theorem 15.2). We will refine the argument to show a more structured lower bound of order $R$ over the class of problems with gap-dependent rate at most $R$.
@@
Next, we claim we can find a partition of $R=\sum_{i=1}^{K-1}R_{i}$ such that $R_{i}\leq\sqrt{(t_{i}-t_{i-1})\cdot(K+1-i)}$ for all $i\in[K-1]$: specifically, let $R_{1}\doteq\min\{\sqrt{(t_{1}-1)\cdot K},R\}$ and recursively define
@@
$\sum_{j:[t_{j-1},t_{j}-1]\cap[s_{1},s_{2}]\neq\emptyset}\left(\frac{1}{2}\sqrt{\frac{K+1-j}{t_{j}-t_{j-1}}}\right)\cdot\frac{(s_{2,j}-s_{1,j})}{K+1-j}\leq(K-2)^{1/2}\cdot\sqrt{\sum_{j:[t_{j-1},t_{j}-1]\cap[s_{1},s_{2}]\neq\emptyset}\frac{s_{2,j}-s_{1,j}}{K+1-j}}.$
@@ C.5   Proof of Theorem 14
We first show (i). Let $a\in\operatorname{argmin}_{a\in[K]}f_{a}(1/T)$ be an optimal arm at round $1$. Let $m\doteq\left\lfloor\beta\right\rfloor$. Then, by Taylor’s Theorem with Lagrange remainder, we have for all $x\in[0,1]$, there exists $\xi\in[0,1]$ such that:
@@
Now, this means arm $a$ must be safe in the sense of Equation 6 for suitable constant $C_{3}$ since the gap at any round cannot exceed $\sqrt{K/T}$. This shows (i).
Next, we show (ii). Fix a real number $\lambda>\sqrt{K/T}$ consider the class $\mathscr{E}$ of environments with

### v2

@@ Appendix B Proof of Dynamic Regret Upper Bound (Theorem 2)
The main idea is that on each significant phase $[\tau_{i},\tau_{i+1},$), any arm $a$ which is at one point optimal, i.e. $\delta_{t}(a)=0$ for some $t\in[\tau_{i},\tau_{i+1})$, must have a gap function $t\mapsto\delta_{t}(a)$ with large variation within the phase since the gap must also at some point be large because of our notion of significant regret (see Fact 11). More generally, if the gap function is $\left\lfloor\beta\right\rfloor$ times differentiable, then we can find an order $\left\lfloor\beta\right\rfloor$ critical point of the gap function across $\left\lfloor\beta\right\rfloor$ different phases using Rolle’s Theorem. Using the definition of Hölder function (Definition 1), we can then bound the derivatives of the gap functions using this critical point. Such bounds can in turn be plugged into an order-$\left\lfloor\beta\right\rfloor$ Taylor approximation of the gap function. Ultimately, these calculations allow us to relate the phase length $\tau_{i+1}-\tau_{i}$ to the smoothness parameters $\beta,\lambda$. The key claim is that each phase must be roughly at least length $T^{\frac{2\beta}{2\beta+1}}\lambda^{-\frac{2}{2\beta+1}}K^{\frac{1}{2\beta+1}}$ which gives the desired regret bound.
@@
$\sum_{i=0}^{{\tilde{L}}}\sqrt{K\cdot(\tau_{i+1}-\tau_{i})}$ $=\sqrt{KT}+\sum_{i:\tau_{i+1}-\tau_{i}>(m+1)\cdot K}\sqrt{K\cdot(\tau_{i+1}-\tau_{i})}$
$\qquad+\sum_{i:\tau_{i+1}-\tau_{i}\leq(m+1)\cdot K}\sqrt{K\cdot(\tau_{i+1}-\tau_{i})},$
@@
Note that, going forward in the proof, we will constrain our attention to significant phases $[\tau_{i},\tau_{i+1})$ such that $\tau_{i+1}\neq T+1$ as the $\sqrt{KT}$ term on the above RHS accounts for this regret contributed by this final phase.
@@
First, suppose $m=0$. Then, we have by Fact 11 that $F(x)=\delta_{x\cdot T}(a)\geq\sqrt{\frac{K}{\tau_{i+1}-\tau_{i}+1}}$ for some $x\in[\tau_{i}/T,\tau_{i+1}/T]$. Combining this with our above display, there exists $x\in[\tau_{i}/T,\tau_{i+1}/T]$ such that for $m=0$:
@@
Then, taking an order-$(m-1)$ Taylor expansion with Lagrange remainder of $F$ about a root $x_{1}\in[\tau_{i}/T,\tau_{i+1}/T)$ (which we already argued exists since $a$ is optimal at some round in $[\tau_{i},\tau_{i+1})$), we have there exists $\xi\in[\tau_{i}/T,\tau_{i+1}/T)$ such that for all $x\in[\tau_{i}/T,\tau_{i+1}/T)$:
$|F(x)|$ $=|F(x)-F(x_{1})|$ ($F(x_{1})=0$)
@@
We next analyze the short significant phases $[\tau_{i},\tau_{i+1})$ where $\tau_{i+1}-\tau_{i}\leq(m+1)\cdot K$. The difficulty here is that we cannot directly apply the same argument as we did for long phases since there may not exist $m+1$ different rounds where an arm is optimal within the phase. To get around this, we will concatenate different short phases together and construct pseudo phases where we can apply the argument as we made for long phases.
@@
we will now further decompose the dynamic regret over short phases using Jensen’s inequality and the fact that each pseudo phase $[n_{j},n_{j+1})$ can contain at most $(m+1)$ short phases $[\tau_{i},\tau_{i+1})$ since all phases are length at least $K$ (Fact 10):
@@
where $J_{1}$ and $J_{2}$ are respectively the number of filled and unfilled pseudo phases. we will first bound $J_{1}$.
@@ C.1   Proof of Theorem 5
$\delta_{\ell}(a)\sum_{t\in P_{\ell}\cap[1,t_{a}-1]}\frac{1}{|\mathcal{S}_{t}|}\leq C_{3}^{2}\frac{\log(T)\sum_{t\in P_{\ell}\cap[1,t_{a}-1]}|\mathcal{S}_{t}|^{-1}}{\delta_{\ell}(a)\sum_{t\in P_{\ell}\cap[1,t_{a}-1]}|\mathcal{S}_{t}|^{-1}}=\frac{C_{3}^{2}\log(T)}{\delta_{\ell}(a)}.$
@@ C.2   Proof of Theorem 6
$\leq C_{3}\sum_{i=1}^{K}\sum_{a\in[K]}\boldsymbol{1}\{a\in\mathcal{S}_{t_{i}-1}\backslash\mathcal{S}_{t_{i}}\}\sqrt{\sum_{t=1}^{t_{i}-1}\frac{\log(T)}{|\mathcal{S}_{t}|}}$ (from Equation 6)
$\leq C_{3}\sqrt{K\sum_{i=1}^{K}\sum_{t=1}^{t_{i}-1}\frac{\log(T)}{|\mathcal{S}_{t}|}}$ (Jensen)
$\leq C_{3}\sqrt{K\sum_{i=1}^{K}|\mathcal{S}_{t_{i-1}}|\sum_{t=t_{i-1}}^{t_{i}}\frac{\log(T)}{|\mathcal{S}_{t_{i-1}}|}}$ (from $\mathcal{S}_{t}\supseteq\mathcal{S}_{t+1}$)
$\leq C_{3}\sqrt{KT\log(T)}.$
@@ C.3   Proof of Theorem 7
$\sum_{s=s_{1}}^{s_{2}}\frac{1}{|\mathcal{A}_{s}|}\geq\max_{a^{\prime}\in[K]}\sum_{s=s_{1}}^{s_{2}}\frac{\hat{\delta}_{s}(a^{\prime},{a^{\sharp}})}{|\mathcal{A}_{s}|}\geq C_{6}\sqrt{\sum_{s=s_{1}}^{s_{2}}\frac{\log(T)}{|\mathcal{A}_{s}|}}\implies\sum_{s=s_{1}}^{s_{2}}\frac{1}{|\mathcal{A}_{s}|}\geq C_{6}^{2}\log(T).$
In particular, this means for $C_{6}>1$ in Algorithm 3: we will have
@@
which violates the definition of the safe arm (Definition 5) for $C_{4}=1$ for $C_{6}$ sufficiently large. Thus, we conclude that ${a^{\sharp}}\in\cap_{t=1}^{T}\mathcal{A}_{t}$.
@@
The first bound is order $\sqrt{\sum_{s=s_{1}}^{s_{2}}\frac{\log(T)}{|\mathcal{A}_{s}|}}$ via the definition of ${a^{\sharp}}$ and the second sum is also the same order by our concentration estimate Equation 14 and eviction criterion (Algorithm 3). This establishes Equation 6 for appropriately large $C_{3}$ (in terms of $C_{6}$). Thus, $\{\hat{t}_{i}\}_{i\in[K]}$ are valid eviction times w.r.t. safe armsets $\{\mathcal{A}_{t}\}_{t=1}^{T}$.
@@ C.4   Proof of Theorem 8
Fix a positive integer $T$, number of arms $K\in[2,T]\cap\mathbb{N}$ and a real number $\Delta\in[0,\sqrt{K/T}]$. Let $\mathscr{E}^{\prime}$ be the class of all environments such that $\sum_{t=1}^{T}\mathbb{E}_{a\sim\operatorname{\operatorname{Unif}}\{[K]\}}[\delta_{t}(a)]\leq R$, and such that $T$ is a valid eviction time w.r.t. initial time $1$ and threshold $C_{3}=1$ in Equation 6. Then, for any algorithm $\pi$, we have:
@@
As in the proof of Theorem 1, we’ll follow a Le Cam’s method style of argument for showing minimax lower bounds in stationary bandits (Lattimore and Szepesvári, 2020, e.g., Theorem 15.2). We will refine the argument to show a more structured lower bound of order $R$ over the class of problems with gap-dependent rate at most $R$.
@@
Next, we claim we can find a partition of $R=\sum_{i=1}^{K-1}R_{i}$ such that
$R_{i}\leq\sqrt{(t_{i}-t_{i-1})\cdot(K+1-i)},$
for all $i\in[K-1]$: specifically, let $R_{1}\doteq\min\{\sqrt{(t_{1}-1)\cdot K},R\}$ and recursively define
@@
$\sum_{j:[t_{j-1},t_{j}-1]\cap[s_{1},s_{2}]\neq\emptyset}\left(\frac{1}{2}\sqrt{\frac{K+1-j}{t_{j}-t_{j-1}}}\right)\cdot\frac{(s_{2,j}-s_{1,j})}{K+1-j}\leq\sqrt{(K-2)\sum_{j:[t_{j-1},t_{j}-1]\cap[s_{1},s_{2}]\neq\emptyset}\frac{s_{2,j}-s_{1,j}}{K+1-j}}.$
@@ C.5   Proof of Theorem 9
We first show 1. Let $a\in\operatorname{argmin}_{a\in[K]}f_{a}(1/T)$ be an optimal arm at round $1$. Let $m\doteq\left\lfloor\beta\right\rfloor$. Then, by Taylor’s Theorem with Lagrange remainder, we have for all $x\in[0,1]$, there exists $\xi\in[0,1]$ such that:
@@
Now, this means arm $a$ must be safe in the sense of Equation 8 for suitable constant $C_{3}$ since the gap at any round cannot exceed $\sqrt{K/T}$. This shows 1.
Next, we show 2. Fix a real number $\lambda>\sqrt{K/T}$ consider the class $\mathscr{E}$ of environments with
@@ Appendix D Auxilliary Lemmas
Appendix E Experimental Results on Synthetic Data
Figure 3: In order, plot of rewards and plot of regret over time $t\in[10^{7}]$.
We present the results of implementing three algorithms for smooth non-stationary bandits: META (Algorithm 1), the Budgeted Exploration (BE) algorithm of Jia et al. (2023), and a random baseline RAND which selects an arm uniformly at random. The code can be found at https://github.com/joesuk/SmoothBandits.
We consider a synthetic environment with trigonometric mean reward functions similar to those used in Jia et al. (2023). In particular, we have $K=2$ arms and trigonometric rewards parametrized by an amplitude $A$, frequency $\nu$, and phase-shift $\varphi$:
$\mu_{1}(t)$ $:=A$
$\mu_{2}(t)$ $:=A-A\sin(2\pi\nu t/T+\varphi),$
where we drew $A=\nu^{-2}$, $\varphi\sim\operatorname{\operatorname{Unif}}\{[0,2\pi]\}$, and $\nu\sim\operatorname{\operatorname{Unif}}\{[2.5,5]\}$. The results of Figure 3 use
$A$ $=0.01444588223139156$
$\nu$ $=8.320088866618766$
$\varphi$ $=1.1478977247810018.$
These reward functions are $C^{\infty}$ smooth in normalized time and are thus Hölder class for any exponent $\beta>0$. In particular, the function $f(x):=A-A\sin(2\pi\nu x+\varphi)$, has Hölder coefficients:
$\lambda_{1}$ $:=\sup_{x\in[0,1]}f^{(1)}(x)=\pi/\nu$
$\lambda_{2}$ $:=\sup_{x\in[0,1]}f^{(2)}(x)=\pi^{2}$
$\lambda_{n}$ $:=\sup_{x\in[0,1]}f^{(n)}(x)=\pi^{n}\cdot\nu^{n-2}.$
In Figure 3, we plot the mean regret curves (with confidence bands of one standard deviation) for $100$ different simulations with a horizon of $T=10^{7}$ and Gaussian reward noise with variance $0.001$. We also plot, in varying shades of blue, the regret curves, for $\beta=1,2,3,\infty$, the theoretical minimax rate of $T^{\frac{\beta+1}{2\beta+1}}\lambda^{\frac{1}{2\beta+1}}K^{\frac{\beta}{2\beta+1}}$ of Theorem 1. Here, “$\beta=\infty$” means we consider the parametric regret rate of $\sqrt{({\tilde{L}}+1)KT}$ in terms of ${\tilde{L}}$ significant shifts, which scales like Equation 2 and lower bounds the minimax regret rate for Hölder class rewards by Theorem 2. For the parameter choices of $A,\nu,\varphi$ above and $T=10^{7}$, one can verify that there are ${\tilde{L}}=4$ significant shifts, and so our “$\beta=\infty$” curve in Figure 3 simply corresponds to the function $t\mapsto\sqrt{10t}$. Note also, for the trigonometric reward model that
$\lim_{n\to\infty}T^{\frac{n+1}{2n+1}}\cdot\lambda_{n}^{\frac{1}{2n+1}}\cdot K^{\frac{n}{2n+1}}=\sqrt{\pi\cdot\nu\cdot KT},$
Thus, this parametric $\sqrt{KT}$ rate does indeed capture the limiting regret as the smoothness $\beta\to\infty$.
The BE algorithm was implemented using the theoretically rate-optimal parameters of Theorem 4.2 of Jia et al. (2023). Our implementation also slightly modifies META (Algorithm 1) to only check the elimination criterion Equation 4 (Lines 2 and 2) over intervals $[t_{0},t]$ of dyadic length, or such that $2^{t-t_{0}+1}=2^{n}$ for some $n\in\mathbb{N}$, which was done to ensure reasonable computation time.
From our plot, we conclude that META indeed outperforms BE (which performs similarly to a naive random baseline here), and is thus able to leverage higher-order smoothness beyond the $\beta=2$ Hölder class accounted for by BE.
Interestingly, we also see that META’s regret curve closely matches the parametric regret rate $\sqrt{({\tilde{L}}+1)KT}$ in terms of significant shifts. This empirically validates the findings of Theorem 2 and Corollary 3.